h(x)=x+1 and g(x)= radical x-6 find (g+h)(6)

Answer by stanbon(75887)

h(x)=x+1 and g(x)= radical x-6 find (g+h)(6)
----
(g+h)(6)
= g(6)+h(6)
= sqrt(6-6) + (6+1)
= 0+7
= 7
=====================
Cheers,
Stan H.